OpenStudy (anonymous):

Solve the system by substitution. 2x+y=-11 3x-4y=11 A.(3, 5) B.(–5, –3) C.(–3, –5) D.(5, 3)

OpenStudy (noseboy908):

first pick what you want so solve for. (get a variable alone on one side)

OpenStudy (anonymous):

2x+y=-11 3x-4y=11 ------------ solve 2x+y=-11 for y

OpenStudy (anonymous):

2x+y=-11 (1) 6+5=11 you have -11 (1) -10-3=-13 not -11 =>A. =>B 3x-4y=11 (1)=>-6-5=-11 A.(3, 5) =>C, (2)=>-9=20=11 B.(–5, –3) C.(–3, –5) (1)=> 10+3=13 you have -11 D.(5, 3) =>D so only C is right?